OpenStudy (anonymous):

John is going to rent a tent to go camping. The cost to rent a camping tent from Woodland Outfitters is a flat rate of $12.00 plus $2.00 for each night the tent is rented. Part a. Let x represent the number of nights John keeps the tent and write a cost function, c(x), for the cost to rent a tent from Woodland Outfitters. Part b. How much will John owe if he uses the tent for three nights?

OpenStudy (anonymous):

This is easy. Look It says that he has a flat rate of $12.00, this will by your y-intercept. So you know you will have a "+ 12" at the end of your function, c(x).

OpenStudy (anonymous):

It says $2 a night, and it says x will equal the number of nights. So that means you will have 2x. So for part A, your function will be c(x) = 2x + 12

OpenStudy (anonymous):

To find out how much John will owe if he uses the tent for three nights, plug in 3 for x in our equation: c(x) = 2x + 12 c(x) = 2(3) + 12 c(x) = 6 + 12 c(x) = 18 So John will owe $18.00 if he uses the tent for 3 days.

OpenStudy (anonymous):

You can give a best answer by clicking "best response". :)

OpenStudy (anonymous):

Thanks man!

OpenStudy (anonymous):

No Problem.

OpenStudy (anonymous):

Just tag me if you need more help.